Trains A and Trains B are traveling in the same directions in parallel tracks. Train A is traveling at 60 miles and hour and Train B is traveling at 64 miles per hour. Train A passes the station at 6:15am. If train B passes the same station at 6:45am, what time will train B catch up to Train A
Answer by checkley77(12844)

D=RT
D=60T
D=64
PROOF:
60*8=64(8-.5)
(T-.5)
SEEING AS THE DISTANCES ARE EQUAL. THEN:
60T=64(T-.5)
60T=64T-32
60T-64T=-32
-4T=-32
T=-32/-4
T=8 HOURS AFTER TRAIN A LEAVES THE STATION THEY WILL MEET.
60*8=64(8-.5)
480=64*7.5
480=480